OpenStudy (kawaiiiimya):

1)An anthropologist finds bone that her instruments measure it as 0.105% of the amount of Carbon-14 the bones would have contained when the person was alive. How long ago did the person die? The half life of carbon 14 is 5,730 years. Round your answer to the nearest thousand. 2) A radio active isotope decays according to the exponential decay equation where t is in days. Round to the thousandths place. For the half life: The half life is the solution (t) of the equation : a/2=ae^−3.524t 3)The Mendes family bought a new house 5 years ago for $100,000. The house is now wor

satellite73 (satellite73):

solve \[\huge \left(\frac{1}{2}\right)^{\frac{t}{5700}}=0.105\]

OpenStudy (kawaiiiimya):

do you want me to solve for T

satellite73 (satellite73):

yeah i made a typo, but you can fix it i am sure

satellite73 (satellite73):

\[\huge \left(\frac{1}{2}\right)^{\frac{t}{5730}}=0.105\]

OpenStudy (kawaiiiimya):

would it be 6?

satellite73 (satellite73):

i really doubt it

satellite73 (satellite73):

first step it change of base formula \[\frac{t}{5730}=\frac{\ln(.105)}{\ln(\frac{1}{2})}\]

OpenStudy (kawaiiiimya):

when I put that into my calculator I get 3.251538767

satellite73 (satellite73):

so then \[t=5730\times \frac{\ln(.105)}{\ln(.5)}\]

OpenStudy (kawaiiiimya):

18631.31713

satellite73 (satellite73):

yeah round a bit maybe

OpenStudy (kawaiiiimya):

to what palce

satellite73 (satellite73):

idk does it say?

OpenStudy (kawaiiiimya):

nearest thousand so 18,631

satellite73 (satellite73):

18631.31713 to the nearest thousand is 1900 to the nearest thousandth is 1863.317

OpenStudy (sainath):

Satellite I need help pls